Calculate the mass of a de Broglie’s particle traveling at a speed of 30m/s
with wavelength 1.49x10-26 nm:
A 0.43 x 10-3 kg.
B 0.15 kg.
C 0.35 x 10-3 kg.
D 0.96 kg.
1
Expert's answer
2020-05-11T20:21:33-0400

"\lambda=\frac{h}{mv}"


"m=\frac{h}{v\lambda}"


"m=\frac{6.63\times10^{-34}}{30\times1.49\times10^{-26}\times10^{-9}}=0.148\times10=1.48kg"